Rocky Guide Service provides guided 1–5 day hiking tours throughout the Rocky Mountains. Wilderness Tours hires Rocky to lead various tours that Wilderness sells. Rocky receives $2,800 per tour day, and shortly after the end of each month Rocky learns whether it will receive a $280 bonus per tour day it guided during the previous month if its service during that month received an average evaluation of "excellent" by Wilderness customers. The $2,800 per day and any bonus due are paid in one lump payment shortly after the end of each month.

On July 1, based on prior experience, Rocky estimated there is a 40% chance it will earn the bonus for July tours. It guided a total of 10 days from July 1–July 15.
On July 16, based on Rocky’s view that it had provided excellent service during the first part of the month, Rocky revised its estimate to an 90% chance it would earn the bonus for July tours. Rocky also guided customers for 15 days from July 16–July 31.
On August 5 Rocky learned it did not receive an average evaluation of “excellent” for its July tours, so it would not receive any bonus for July, and received all payment due for the July tours.

Rocky bases estimates of variable consideration on the expected value it expects to receive.
Answer #1

Expected value means whatever the chances of receipt, it must be accounted in the books.

No.

Date

Journal entry

Debit

Credit

1

July 15

Accounts Receivable (2800 * 10 days)

$ 28,000

Bonus receivable (2800 * 10 * 0.4)

$ 1120

        Service Revenue

$ 29,120

( To record revenue from tour given from July 1-15)

2

July 31

Accounts Receivable (2800*15 days)

$ 42,000

Bonus receivable (280 * 25 days * 0.9) -1120

$ 5,180

        Service Revenue

$ 47,180

( To record revenue from tour given from July 16-31)

3

Aug 5

Cash (42,000 + 18,000)

$ 70,000

     Accounts Receivable

$ 70,000

(To record receipt of payment)

4

Aug 5

Service revenue (1120 + 5180)

$ 7,000

        Bonus receivable

$ 7,000

(To record adjustment to revenue as bonus won’t be received later)

Explanations:

  1. For the period July 1-15, Rocky estimates 40% chance that bonus will be earned. Using expected amount approach, Rocky should record 40% bonus for the period.
  2. For period July 16-31, the revenue earned = 2800*15 = $ 42,000. Further, Rocky estimates 90% chance that bonus will be received. Using expected approach, bonus estimate would be 280 * (10+15 days) = $ 7,000 less 1120 which is already recognized = $ 5180.
Answer #2

1 - there should be no accrual for "most likely" 40% accrual since its below 50%.

2 - sicne the most likely is above 50% you need to accrue the entire amoutn of potential bonus and not the 90% of it.

answered by: olessia silakova
Answer #3

image.png

answered by: olessia silakova